Calculus II, Spring 2019 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 14

1. Section 15.7: Problems 9, 11, 13, 14, 15, 19, 21, 31, 39, 43, 45.

2. Section 15.8: Problems 5, 7, 9, 15, 23.

3. Chap 15, additional and advanced exercises: Problems 11, 12.